Hoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singapore. The company has determined that if a truck is driven 105,000 kilometers during a year, the average operating cost is 11.6 cents per kilometer. If a truck is driven only 70,000 kilometers during a year, the average operating cost increases to 13.3 cents per kilometer.

Required:

1. Using the high-low method, estimate the variable operating cost per kilometer and the annual fixed operating cost associated with the fleet of trucks.

2. Express the variable and fixed costs in the form Y = a + bX.

3. If a truck were driven 87,500 kilometers during a year, what total operating cost would you expect to be incurred?
Answer #1

Total cost:

105,000 kilometres = 105,000 X $11.6 = $1,218,000

70,000 kilometres = 70,000 X $13.3 = $931,000

1.

Variable cost per kilometre = ($1,218,000 - $931,000) / (105,000 - 70,000) = $8.20

Fixed cost = $1,218,000 - (105,000 X $8.20) = $357,000

2.

Y = 357,000 + b8.20

3.

Total operating cost = $357,000 + (87,500 X $8.20) = $1,074,500